OpenStudy (joshsellsguns):

3x + 5 <= 2x- 8 solve the inequality this is the example they gave: 3x + 5 > -1 3x > 6 x > -2

OpenStudy (anonymous):

You need to find x and put it to one side. To do this, you need to do opposite factors (e.g plus will rid of negative, multiply will rid of divide and vice verse). Imagine a scale - you need to do both sides or it won't be equal. If you want, I can guide you through this.

OpenStudy (joshsellsguns):

Yes I would like that. I am kind of getting the general idea but I could use alittle more guidance on it.

OpenStudy (anonymous):

3x + 5 <= 2x- 8 I'm going to get rid of 5, to leave 3x. So, -5 3x +5 - 5 <= 2x - 8 -5 3x <= 2x - 13 Keep going until you cannot simplify it anymore. Does that make sense?

OpenStudy (anonymous):

I'll be leaving soon, so best if you respond quickly.

OpenStudy (joshsellsguns):

Okay sorry. And yes that makes sense. Thank you.

OpenStudy (anonymous):

No problem, is there anymore questions or you understand how to do this? By the way, welcome to openstudy :)

OpenStudy (joshsellsguns):

Tank you and I got it now.

OpenStudy (joshsellsguns):

*Thank

OpenStudy (anonymous):

Good on you, good luck on your future questions! ^_^/